### v2

@@ 6 The Dutch-book theorem for complete consistency
The Dutch-book theorem for complete consistency generalizes the insights from the example: incongruities among generalized odds ratios can be exploited for Dutch-booking and can only be displayed by a belief system that is not completely consistent. The theorem also shows the opposite implications: when there is no incongruity between generalized odds ratios, the agent’s belief system is completely consistent and he cannot be Dutch-booked. Thus, coherence among generalized odds ratios fully characterizes complete consistency and impossibility of Dutch-booking.
@@
Fix a belief system $\mu$. The following are equivalent.
@@
$\mathbf{1\Rightarrow 2)}$
Note preliminarly that, by Condition 1, if a state $s$ has a DOR with another state $s^{\prime}$, all the DORs between $s$ and $s^{\prime}$ take the same value; [^12] Recall that an “indeterminate DOR” betwen $s$ and $s^{\prime}$ is just not considered a DOR, and likewise for GOR. We say that $s$ has a DOR with $s^{\prime}$ when there exists $h\in\mathcal{H}$ such that $s,s^{\prime}\in S(h)$ and $\mu(s|h)+\mu(s^{\prime}|h)>0$, the DOR between $s$ and $s^{\prime}$ being $\mu(s|h)/\mu(s^{\prime}|h)$. Similarly, we say that $s$ has a GOR with $s^{\prime}$ when it is possible to compute at least one GOR between $s$ and $s^{\prime}$. call it $o(s,s^{\prime})$.
Let $P^{1}$ be the set of all $s\in\mathcal{S}$ that have no zero GOR with any $s^{\prime}$.
Let $P^{2}$ be the set of all $s\in\mathcal{S}\backslash P^{1}$ that have no zero GOR with any $s^{\prime}\in\mathcal{S}\backslash P^{1}$.
Let $P^{3}$ be the set of all $s\in\mathcal{S}\backslash\left(P^{1}\cup P^{2}\right)$ that have no zero GOR with any $s^{\prime}\in\mathcal{S}\backslash\left(P^{1}\cup P^{2}\right)$.
@@
Claim 1: there exists $n\geq 1$ such that $(P^{1},...,P^{n})$ is a partition of $\mathcal{S}$, with non-empty elements.
@@
Proof: Let $n$ be the smallest non-negative integer such that $P^{n+1}=\emptyset$; it exists because $\mathcal{S}$ is finite and the sets are disjoint by construction. There remains to show that $\cap_{m=1}^{n}P^{m}=\mathcal{S}$ (which also implies that $n\geq 1$). Let $\widetilde{S}=\mathcal{S}\backslash\left(\cap_{m=1}^{n}P^{m}\right)$ and suppose by contraposition that $\widetilde{S}\not=\emptyset$. The fact that $P^{n+1}=\emptyset$ implies that every $s\in\widetilde{S}$ has a zero GOR with some $s^{\prime}\in\widetilde{S}$. Consider a directed graph with nodes $\widetilde{S}$ and arrows that represent a zero GOR between the start and the end node. Since every node is the start of an arrow, the graph has a cycle. So, the product of all the (zero) GORs along the cycle is a zero GOR between $s^{1}$ and itself. Inverting all the DORs in this GOR, we obtain an infinite GOR between $s^{1}$ and itself. This violates the assumption that all the GORs between $s^{1}$ and itself are identical by Condition 1. $\square$
@@
Claim 2: there exists a LCPS $\bar{\mu}=(\nu^{1},...,\nu^{n})$ such that, for each $m=1,...,n$, (a) $\mathrm{supp}\nu^{m}=P^{m}$ and (b), for every pair $\left(s,s^{\prime}\right)\in P^{m}\times P^{m}$ which has a DOR $o(s^{\prime},s)$, $\nu^{m}(s^{\prime})/\nu^{m}(s)=o(s^{\prime},s)$.
@@
Proof: For each $m=1,...,n$, we will construct a probability measure $\nu^{m}$ that satisfies (a) and (b); then, $(\nu^{1},...,\nu^{n})$ is an LCPS because, by Claim 1, $(P^{1},...,P^{n})$ partitions $\mathcal{S}$.
@@
Fix $m=1,...,n$ and construct a graph with nodes $P^{m}$ and connections between nodes that have a DOR. For every connected component $\widetilde{P}\subseteq P^{m}$, pick a state $s$ arbitrarily, let $\widetilde{\nu}(s)=1$, and progressively let $\widetilde{\nu}(s^{\prime\prime})=\widetilde{\nu}(s^{\prime})\cdot o(s^{\prime\prime},s^{\prime})$ for every $s^{\prime\prime}\in\widetilde{P}$, where $o(s^{\prime\prime},s^{\prime})$ is a DOR between $s^{\prime\prime}$ and some $s^{\prime}\in\widetilde{P}$ for which $\widetilde{\nu}(s^{\prime})$ was already defined; by definition of $P^{m}$, $o(s^{\prime\prime},s^{\prime})$ is not zero and not infinite (otherwise $o(s^{\prime},s^{\prime\prime})$ would be zero). So, we can obtain a probability measure $\nu^{m}$ with support $P^{m}$ by rescaling $\widetilde{\nu}$.
@@
To show that $\nu^{m}$ satisfies the desiderata, fix $s,s^{\prime}\in P^{m}$ that have a DOR $o(s^{\prime},s)$, thus $s$ and $s^{\prime}$ are in the same connected component. Then, the graph has a path $(\bar{s}^{1},...,\bar{s}^{T})$ with $\bar{s}^{1}=s$ and $\bar{s}^{T}=s^{\prime\prime}$ such that, for each $t=2,...,T$, $\widetilde{\nu}(\bar{s}^{t})=\widetilde{\nu}(\bar{s}^{t-1})\cdot o(\bar{s}^{t},\bar{s}^{t-1})$, where $o(\bar{s}^{t},\bar{s}^{t-1})\in\mathbb{R}^{++}$ is a DOR between $\bar{s}^{t}$ and $\bar{s}^{t-1}$. (Note that, if $\widetilde{\nu}(\bar{s}^{t-1})$ was defined as $\widetilde{\nu}(\bar{s}^{t})\cdot o(\bar{s}^{t-1},\bar{s}^{t})$ and not vice versa, there is also a DOR $o(\bar{s}^{t},\bar{s}^{t-1})=1/o(\bar{s}^{t-1},\bar{s}^{t})$ between $\bar{s}^{t}$ and $\bar{s}^{t-1}$.) Therefore
@@
First we show that $\mathrm{supp}\mu(\cdot|h)=P^{m}\cap S(h)$.
Fix $s\in\mathrm{supp}\mu(\cdot|h)$. Thus, $s\in S(h)$. Hence, for every $k<m$, $s\not\in P^{k}$, as $\mathrm{supp}\nu^{k}=P^{k}$ but $\nu^{k}(S(h))=0$. Now we show that, for every $k>m$, $s\not\in P^{k}$ either, so that $s\in P^{m}$. Suppose by contradiction that $s\in P^{k}$. As $\mathrm{supp}\nu^{m}=P^{m}$ and $\nu^{m}(S(h))>0$, there exists $\bar{s}\in S(h)\cap P^{m}$ such that $\nu^{m}(\bar{s})>0$. Since $s\in\mathrm{supp}\mu(\cdot|h)$, $s$ has a non-zero DOR with every $s^{\prime}\in S(h)$, so in particular with $\bar{s}$. Since $s\not\in P^{1}\cup...\cup P^{m}$, $s$ has a zero GOR with some $s^{\prime}\in P^{m}\cup...\cup P^{n}$. The two things combined impy that $\bar{s}$ has a zero GOR with $s^{\prime}$ too. But this contradicts $\bar{s}\in P^{m}$.
Now fix $s\not\in\mathrm{supp}\mu(\cdot|h)$. If $s\not\in S(h)$, we are done. If $s\in S(h)$, $s\not\in P^{m}$, because $s$ has a zero DOR with every $s^{\prime}\in\mathrm{supp}\mu(\cdot|h)$ and we have just shown that $\mathrm{supp}\mu(\cdot|h)\subseteq P^{m}$.
@@
$\mathbf{2\Rightarrow 3)}$
Let $\bar{\mu}=(\mu^{1},...,\mu^{n})$ be a LCPS from which the belief system is derived. For each $m=1,...,n$, let $S^{m}$ denote the support of $\mu^{m}$, and let $H^{m}$denote the set of all $h\in\mathcal{H}$ such that $\mu(\cdot|h)$ is derived from $\mu^{m}$.
Fix a system of gambles $g=(g(\cdot|h))_{h\in\mathcal{H}}$. Suppose that $i$ is willing to accept $g$. We show that $g$ is not a Dutch book. Suppose that $g$satisfies condition (1) of a Dutch book, so that,
@@
The proof is by induction on $m$. Fix $m=1,...,n$ and suppose by way of induction that for every $k<m$, $g(s|h)=0$ for all $s\in\mathcal{S}$ and $h\in H^{k}$. (This is vacuously true for $m=1$.) Fix $s\in S^{m}$. For every $k>m$ and $h\in H^{k}$, $p(h|s)=0$, otherwise $\mu(\cdot|h)$ could be derived from $\mu^{m}$. By this and the induction hypothesis, inequality (5) yields
@@
Since $\sum_{s^{\prime}}p(h|s^{\prime})\mu^{m}(s^{\prime})>0$ for every $h\in H^{m}$, we can rewrite (6) as
$\sum_{h\in H^{m}}\sum_{s\in\mathcal{S}}\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}}p(h|s^{\prime})\mu^{m}(s^{\prime})}g(s|h)\leq 0.$ (7)
@@
Since $\mu(\cdot|h)$ can be derived from $\mu^{m}$ with Bayes rule, we can rewrite (8) as
$\sum_{s\in\mathcal{S}}\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}}p(h|s^{\prime})\mu^{m}(s^{\prime})}g(s|h)\geq 0.$ (9)
@@
$\forall h\in H^{m},\text{ \ \ }\sum_{s\in S}\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}}p(h|s^{\prime})\mu^{m}(s^{\prime})}g(s|h)=0,$

### v3

@@ 6 The Dutch-book theorem for complete consistency
The Dutch-book theorem for complete consistency generalizes the insights from the example: incongruities among generalized odds ratios can be exploited for Dutch-booking and can only be displayed by a belief system that is not completely consistent. The theorem also shows the opposite implications: when there is no incongruity between generalized odds ratios, the agent’s belief system is completely consistent, and he cannot be Dutch-booked. Thus, coherence among generalized odds ratios fully characterizes complete consistency and impossibility of Dutch-booking.
@@
Fix a belief system $\mu$. The following are equivalent:
@@
Before proving the theorem, we conclude the running example.
Example 5
Suppose that Larry has the belief system $\hat{\mu}$ constructed before. As we observed, $\hat{\mu}$ is not completely coherent. Thus, by Theorem 2, $\hat{\mu}$ is not completely consistent, and Larry can be Dutch-booked. In particular, he is willing to accept Dutch book $\hat{g}$:
$\sum_{s\in\mathcal{S}}\hat{\mu}(s|h)\hat{g}(s|h)=\frac{3}{4}\cdot 9+\frac{1}{4}(-10)+0\cdot 0>0\qquad\forall h\in\left\{sm,mp,ps\right\}.$
Thus, no matter the fireworks’ location, Larry ends up with equal probability in each of the other two locations, and hence he suffers an expected loss from this betting behavior. The lack of complete consistency means that Larry’s beliefs cannot be derived from a (lexicographic) prior under the true DGP. This could be because Larry does not have a prior; after all, he does not need one, as he cannot choose which bus to take. Or, it could be that Larry updates a prior with a misspecified DGP, perceiving a correlation between bus arrivals and the state when, in fact, there is none. In Section 7, we will instead provide opposite examples of correlation neglect. $\triangle$
@@
$\mathbf{1\Rightarrow 2)}$ Note preliminarily that, by Condition 1, if a state $s$ has a DOR with another state $s^{\prime}$, all the DORs between $s$ and $s^{\prime}$ take the same value; [^15] Recall that by definition an “indeterminate DOR” between $s$ and $s^{\prime}$ is just not considered a DOR, and likewise for GOR. We say that $s$ has a DOR with $s^{\prime}$ when there exists $h\in\mathcal{H}$ such that $s,s^{\prime}\in S(h)$ and $\mu(s|h)+\mu(s^{\prime}|h)>0$, the DOR between $s$ and $s^{\prime}$ being $\left(\mu(s|h)p(h|s^{\prime})\right)/\left(\mu(s^{\prime}|h)p(h|s)\right)$. Similarly, we say that $s$ has a GOR with $s^{\prime}$ when it is possible to compute at least one GOR between $s$ and $s^{\prime}$. call it $o(s,s^{\prime})$.
Let $P^{1}$ be the set of all $s\in\mathcal{S}$ such that for all $s^{\prime}\in\mathcal{S}$, all the GORs between $s$ and $s^{\prime}$ are different from 0.
Let $P^{2}$ be the set of all $s\in\mathcal{S}\backslash P^{1}$ such that for all $s^{\prime}\in\mathcal{S}\backslash P^{1}$, all the GORs between $s$ and $s^{\prime}$ are different from 0.
Let $P^{3}$ be the set of all $s\in\mathcal{S}\backslash\left(P^{1}\cup P^{2}\right)$ such that for all $s^{\prime}\in\mathcal{S}\backslash\left(P^{1}\cup P^{2}\right)$, all the GORs between $s$ and $s^{\prime}$ are different from 0.
@@
There exists $n\geq 1$ such that $(P^{1},...,P^{n})$ is a partition of $\mathcal{S}$, with non-empty elements.
@@
Proof of the Claim. Let $n$ be the smallest non-negative integer such that $P^{n+1}=\emptyset$; it exists because $\mathcal{S}$ is finite and the sets are disjoint by construction. There remains to show that $\cup_{m=1}^{n}P^{m}=\mathcal{S}$ (which also implies that $n\geq 1$). Let $\widetilde{S}=\mathcal{S}\backslash\left(\cup_{m=1}^{n}P^{m}\right)$ and suppose by contraposition that $\widetilde{S}\not$$\emptyset$. The fact that $P^{n+1}=\emptyset$ implies that every $s\in\widetilde{S}$ has a zero GOR with some $s^{\prime}\in\widetilde{S}$. Consider a directed graph with nodes $\widetilde{S}$ and arrows that represent a zero GOR between the start and the end node. Since every node is the start of an arrow, the graph has a cycle. So, the product of all the (zero) GORs along the cycle is a zero GOR between some $s^{1}\in\widetilde{S}$ and itself. Inverting all the DORs in this GOR, we obtain an infinite GOR between $s^{1}$ and itself. This violates the assumption that all the GORs between $s^{1}$ and itself are identical by Condition 1. $\square$
Claim 2
@@
There exists a LCPS $\bar{\mu}=(\nu^{1},...,\nu^{n})$ such that, for each $m\in\left\{1,...,n\right\}$, (a) $\mathrm{supp}\nu^{m}=P^{m}$ and (b), for every pair $\left(s,s^{\prime}\right)\in P^{m}\times P^{m}$ which has a DOR $o(s^{\prime},s)$, $\nu^{m}(s^{\prime})/\nu^{m}(s)=o(s^{\prime},s)$.
@@
Proof of the Claim. For each $m\in\left\{1,...,n\right\}$, we will construct a probability measure $\nu^{m}$ that satisfies (a) and (b); then, $(\nu^{1},...,\nu^{n})$ is an LCPS because, by Claim 1, $(P^{1},...,P^{n})$ partitions $\mathcal{S}$.
@@
Fix $m\in\left\{1,...,n\right\}$ and construct a graph with nodes $P^{m}$ and connections between nodes that have a DOR. For every connected component $\widetilde{P}\subseteq P^{m}$, pick a state $s$ arbitrarily, let $\widetilde{\nu}(s)=1$, and progressively let $\widetilde{\nu}(s^{\prime\prime})=\widetilde{\nu}(s^{\prime})\cdot o(s^{\prime\prime},s^{\prime})$ for every $s^{\prime\prime}\in\widetilde{P}$, where $o(s^{\prime\prime},s^{\prime})$ is a DOR between $s^{\prime\prime}$ and some $s^{\prime}\in\widetilde{P}$ for which $\widetilde{\nu}(s^{\prime})$ was already defined; by definition of $P^{m}$, $o(s^{\prime\prime},s^{\prime})$ is not zero and not infinite (otherwise $o(s^{\prime},s^{\prime\prime})$ would be zero). So, we can obtain a probability measure $\nu^{m}$ with support $P^{m}$ by rescaling $\widetilde{\nu}$.
@@
To show that $\nu^{m}$ satisfies the desiderata, fix $s,s^{\prime}\in P^{m}$ that have a DOR $o(s^{\prime},s)$, thus $s$ and $s^{\prime}$ are in the same connected component. Then, the graph has a path $(\bar{s}^{1},...,\bar{s}^{T})$ with $\bar{s}^{1}=s$ and $\bar{s}^{T}=s^{\prime}$ such that, for each $t\in\left\{2,...,T\right\}$, $\widetilde{\nu}(\bar{s}^{t})=\widetilde{\nu}(\bar{s}^{t-1})\cdot o(\bar{s}^{t},\bar{s}^{t-1})$, where $o(\bar{s}^{t},\bar{s}^{t-1})\in\mathbb{R}^{++}$ is a DOR between $\bar{s}^{t}$ and $\bar{s}^{t-1}$. That is, there is a path in the graph such that for every two consecutive states, the value of $\tilde{\nu}$ for one state has been defined by the other. (Note that, if $\widetilde{\nu}(\bar{s}^{t-1})$ was defined as $\widetilde{\nu}(\bar{s}^{t})\cdot o(\bar{s}^{t-1},\bar{s}^{t})$ and not vice versa, there is also a DOR $o(\bar{s}^{t},\bar{s}^{t-1})=1/o(\bar{s}^{t-1},\bar{s}^{t})$ between $\bar{s}^{t}$ and $\bar{s}^{t-1}$.) Therefore
@@
First, we show that $\mathrm{supp}\mu(\cdot|h)=P^{m}\cap S(h)$.
Fix $s\in\mathrm{supp}\mu(\cdot|h)$. Thus, $s\in S(h)$. Hence, for every $k<m$, $s\not\in P^{k}$, as $\mathrm{supp}\nu^{k}=P^{k}$ but $\nu^{k}(S(h))=0$. Now we show that, for every $k>m$, $s\not\in P^{k}$ either, so that $s\in P^{m}$. Suppose by contradiction that $s\in P^{k}$. As $\mathrm{supp}\nu^{m}=P^{m}$ and $\nu^{m}(S(h))>0$, there exists $\bar{s}\in S(h)\cap P^{m}$ such that $\nu^{m}(\bar{s})>0$. Since $\in\mathrm{supp}\mu(\cdot|h), ss$ has a non-zero DOR (which can be finite or infinite) with every $s^{\prime}\in S(h)$, so in particular with $\bar{s}$. Since $\not\in\cup_{i=1}^{m}P^{i}, ss$ has a zero GOR with some $s^{\prime}\in\cup_{i=m}^{n}P^{i}$. The two things combined imply that $\bar{s}$ has a zero GOR with $s^{\prime}$ as well. But this contradicts $\bar{s}\in P^{m}$.
Now fix $s\not\in\mathrm{supp}\mu(\cdot|h)$. If $s\not\in S(h)$, we are done. If $s\in S(h)$, $s\not\in P^{m}$, because $s$ has a zero DOR with every $s^{\prime}\in\mathrm{supp}\mu(\cdot|h)$, and we have just shown that $\mathrm{supp}\mu(\cdot|h)\subseteq P^{m}$.
@@
$\mathbf{2\Rightarrow 3)}$ Let $\bar{\mu}=(\mu^{1},...,\mu^{n})$ be an LCPS from which the belief system is derived. For each $m\in\left\{1,...,n\right\}$, let $S^{m}$ denote the support of $\mu^{m}$, and let $H^{m}\$denote the set of all $h\in\mathcal{H}$ such that $\mu(\cdot|h)$ is derived from $\mu^{m}$.
Fix a system of gambles $g=(g(\cdot|h))_{h\in\mathcal{H}}$. Suppose that $i$ is willing to accept $g$. We show that $g$ is not a Dutch book. Suppose that gg\ satisfies condition (1) of a Dutch book, so that,
@@
The proof is by induction on $m$. Fix $m\in\left\{1,...,n\right\}$ and suppose by induction that for every $k<m$, $g(s|h)=0$ for all $s\in\mathcal{S}$ and $h\in H^{k}$. (This is vacuously true for $m=1$.) Fix $s\in S^{m}$. For every $k>m$ and $h\in H^{k}$, $p(h|s)=0$, otherwise $\mu(\cdot|h)$ could be derived from $\mu^{m}$. By this and the induction hypothesis, inequality (5) yields
@@
Since $\sum_{s^{\prime}\in\mathcal{S}}p(h|s^{\prime})\mu^{m}(s^{\prime})>0$ for every $h\in H^{m}$, we can rewrite (6) as
$\sum_{h\in H^{m}}\sum_{s\in\mathcal{S}}\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}\in\mathcal{S}}p(h|s^{\prime})\mu^{m}(s^{\prime})}g(s|h)\leq 0.$ (7)
@@
Since $\mu(\cdot|h)$ can be derived from $\mu^{m}$ by Bayes rule, we can rewrite (8) as
$\forall h\in H^{m},\text{ \ \ }\sum_{s\in\mathcal{S}}\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}\in\mathcal{S}}p(h|s^{\prime})\mu^{m}(s^{\prime})}g(s|h)\geq 0.$ (9)
@@
$\forall h\in H^{m},\text{ \ \ }\sum_{s\in\mathcal{S}}\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}\in\mathcal{S}}p(h|s^{\prime})\mu^{m}(s^{\prime})}g(s|h)=0,$